A car drives over a hilltop that has a radius of curvature 0.114 km at the top of the hill. At what speed would the car be traveling when it tires just barely lose contact with the road when the car is at the top of the hill? (Give answer to the nearest 0.1 m/s)
Answer #1

car is in circular motion having centripetal acceleration procided by mg

so in earth's reference, so mv^2 /Roc is the force outward that car is feeling

balancing condition here is mg = mv^2/R

or

v^2 = gR

V = sqrt(g R)

v=sqrt(9.81*114)

V = 33.44 m/s^2